Two coils A and B are connected in series across a 110V, 60Hz supply. The resistance of coil A is 2  and inductance of coil B is 0.004H. If the input from the supply is 2.4kW and 1.8kVAR, find the inductance of coil A and resistance of coil B. Also calculate the voltage across each coil.


1
Expert's answer
2021-02-22T10:19:21-0500

Voltage applied, "V = 110 V"

Frequency, "f = 60 Hz"

The resistance of coil A, "R_A = 2 \Omega"

Then it's inductance, "R_A = 2\pi f L_A \implies L_A = \frac{R_A}{2\pi f} = 0.0053 H"


Inductance of coil B, "L_B = 0.004H"

Resistance of coil B, "R_B = 2\pi fL_B = (2\times \pi \times 60 \times0.004) = 1.5 \Omega"


Since these are in series, so current in A and B will be same. Current in the circuit is given by, "I = \frac{V}{R} = \frac{V}{R_A+R_B}=\frac{110}{2+1.5} = \frac{110}{3.5} A"


Voltage across A, "V_A = IR_A = \frac{110}{3.5}\times 2 = 62.86 V"


Voltage across B, "V_B=IR_B =\frac{110}{3.5}\times 1.5 = 47.14V"